# On relative tractor bundles

### v1

@@ 3.4. Example: Legendrean contact structures
Legendrean contact structures (also called Lagrangean contact structures following the article [Takeuchi] in which they were first studied) are an example of parabolic contact structures, so they can be viewed as a refinement of contact structures. Explicitly, a Legendrean contact structure on a smooth manifold of odd dimension $12n+$ 1 is given by a contact structure $H\subset TM$ together with a decomposition $H=E\oplus F$ of $H$ as a direct sum of two subbundles of rank $n$, which are Legendrean in the sense that the Lie bracket of two sections of one of the subbundles always is a section of $H$. Otherwise put, for any $x\in M$ the fibers $E_{x},F_{x}\subset H_{x}\subset T_{x}M$ are maximal isotropic subspaces with respect to the non-degenerate bilinear form $\mathcal{L}_{x}:H_{x}\times H_{x}\to T_{x}M/H_{x}$ induced by the Lie bracket of vector fields. The interest in these structures mainly comes from their relation to the geometric theory of differential equations, see e.g. [MFMZ] and to CR geometry, see e.g. [Doubrov-Merker-The]. Indeed, Legendrean contact structures and partially integrable almost CR structures are two real forms of the same complex geometric structure.
@@
The relation to parabolic geometries comes from the homogeneous model of such geometries, which is the partial flag manifold $F_{1,n+1}(\mathbb{R}^{n+2})$ of lines contained in hyperplanes in $\mathbb{R}^{n+2}$. This is a homogeneous space of the group $G:=PGL(n+2,\mathbb{R})$ and the stabilizer of the standard flag $\mathbb{R}\subset\mathbb{R}^{n+1}\subset\mathbb{R}^{n+2}$ descends to a parabolic subgroup $Q\subset G$. Indeed, $Q=P\cap\tilde{P}$ for two maximal parabolic subgroups of $G$, the stabilizers of $\mathbb{R}\subset\mathbb{R}^{n+2}$ and $\mathbb{R}^{n+1}\subset\mathbb{R}^{n+2}$. Correspondingly, there is a canonical homogeneous projection $G/Q\to G/P=\mathbb{R}P^{n+1}$ and it is easy to see that this identifies $G/Q$ with the projectivized cotangent bundle $\mathcal{P}(T^{*}\mathbb{R}P^{n+1})$. This gives rise to a canonical contact structure on $G/Q$ for which the vertical subbundle is well known to be Legendrean. The second Legendrean subbundle making $G/Q$ into a Legendrean contact manifold is the vertical bundle of the canonical projection $G/Q\to G/\tilde{P}\cong\mathbb{R}P^{(n+1)*}$.
@@
(3.2) $\begin{pmatrix}\mathfrak{g}_{(0,0)}&\mathfrak{g}_{(1,0)}&\mathfrak{g}_{(1,1)}\\ \mathfrak{g}_{(-1,0)}&\mathfrak{g}_{(0,0)}&\mathfrak{g}_{(0,1)}\\ \mathfrak{g}_{(-1,-1)}&\mathfrak{g}_{(0,-1)}&\mathfrak{g}_{(0,0)}\end{pmatrix}$
@@
In particular, we just get one representation $\mathbb{V}_{i^{\prime}}$ in this case, namely $\mathbb{V}_{-1}=\mathfrak{g}/\mathfrak{p}$ and the $Q$-invariant filtration of this representation consists of the single invariant subspace $(\mathfrak{g}_{(-1,0)}\oplus\mathfrak{p})/\mathfrak{p}$. The corresponding bundles are easy to interpret, too
@@
General results on parabolic geometries imply that for any Legendrean contact structure $(M,H=E\oplus F)$ there is a normal Cartan geometry $(p:\mathcal{G}\to M,\omega)$ of type $(G,Q)$ which is unique up to isomorphism, see Section 4.2.3 of [book]. It is a general fact about Cartan geometries that $\mathcal{G}\times_{Q}(\mathfrak{g}/\mathfrak{q})\cong TM$ and the characterizing property of the geometry (apart from normality) is that the $Q$-invariant subspaces $(\mathfrak{g}_{(-1,0)}\oplus\mathfrak{q})/\mathfrak{q}\subset\mathfrak{g}/\mathfrak{q}$ and $(\mathfrak{g}_{(0,-1)}\oplus\mathfrak{q})/\mathfrak{q}\subset\mathfrak{g}/\mathfrak{q}$ correspond to the subbundles $E$ and $F$, respectively. Since $\mathfrak{g}_{(0,-1)}\oplus\mathfrak{q}=\mathfrak{p}$ we conclude that $T_{\rho}M=F$ in this case, and that our basic tractor bundle $\mathcal{V}_{-1}M$ is simply $TM/F$. The natural filtration of this bundle is given by the single invariant subspace $H/F\subset TM/F$.
@@
Now we can collect our results in the case of Legendrean contact structures. We restrict to the case $n>1$, i.e. $\dim(M)\geq 5$. The three-dimensional case can be easily dealt with directly. It is not very interesting from the point of view of relative BGG theory, however, since relative BGG sequences consist only of a single operator in this dimension.
@@
In the current situation, relative tractor bundles correspond to representations of the general linear group $GL(n+1,\mathbb{R})$, so they correspond to natural vector bundles on smooth manifolds of dimension $n+1$. The relative tractor bundle $\mathcal{V}M=TM/F$ corresponds to the standard representation of $GL(n+1,\mathbb{R})$ on $\mathbb{R}^{n+1}$, so via tensorial constructions as in §3.3 we obtain all analogs of tensor bundles. It is clear that the latter are by far the most important natural vector bundles (in fact, we are not aware of interesting applications of other types of natural bundles). Thus it should be clear that we obtain (up to isomorphism) all the important relative tractor bundles on Legendrean contact structures via the constructions of §3.3, but it is difficult to make a precise statement in this direction. Instead of trying to do this, we just prove a result on the resulting representations in the following theorem, which shows the richness of the class.
Theorem 10.
@@
Let $(M,H=E\oplus F)$ be a Legendrean contact structure of dimension $2n+1\geq 5$ such that the distribution $F\subset TM$ is involutive and consider the relative tractor bundle $\mathcal{V}M:=TM/F$. Applying tensorial constructions to $\mathcal{V}M$ as in §2.3 we arrive at a class of relative tractor bundles that correspond to representations of $P/P_{+}\cong GL(n+1,\mathbb{R})$ which contains all one-dimensional representations and whose restrictions to $\mathfrak{sl}(n+1,\mathbb{R})\subset\mathfrak{p}/\mathfrak{p}_{+}$ exhaust all finite dimensional representations of this Lie algebra.
@@
For any of these relative tractor bundles and any local leaf space $\psi:U\to N$ for the distribution $F$ with connected fibers, the induced relative BGG sequence defines a fine resolution of the sheaf of pullbacks of sections of a tensor bundle on $N$. This tensor bundle is obtained by applying the tensorial constructions from above to the tangent bundle $TN$.
@@
We first need some information on the curvature of Legendrean contact structures that is available in the literature. For $n>1$, Proposition 4.3.2 of [book] completely describes the harmonic curvature of Legendrean contact structures: There are two harmonic curvature quantities of homogeneity one, which are sections of the bundles $\Lambda^{2}E^{*}\otimes F$ and $\Lambda^{2}F^{*}\otimes E$, respectively. These turn out to be induced by the Lie bracket of vector fields and they are exactly the obstructions to involutivity of the distributions $E$ and $F$. So in particular, one of these components is exactly the obstruction to involutivity of $T_{\rho}M=F$. The third and last harmonic curvature quantity has homogeneity two and can be interpreted as a section of $E^{*}\otimes F^{*}\otimes L(E,E)$ (with some symmetries).
@@
Hence involutivity of $F$ implies that the harmonic curvature vanishes upon insertion of two elements of $F$ while the components of torsion type even vanish upon insertion of one element of $F$. These are exactly the assumptions needed to apply part (1) of Proposition 4.18 of [Rel-BGG2]. The proof of this proposition then shows that the full Cartan curvature has the same properties, which in our language means that $\tau(T_{\rho}M,TM)=0$, so the assumptions of Corollary 9 are satisfied. On the other hand, Proposition 4.18 of [Rel-BGG2] says that the assumptions of part (1) of Theorem 4.11 of [Rel-BGG2] are satisfied. This shows that any relative BGG sequence is a fine resolution of the sheaf of local parallel sections for the relative tractor connection. Together with Corollary 9, this implies all claimed properties of the relative BGG sequence for $\mathcal{V}M=TM/F$. For bundles obtained via tensorial constructions, they then follow readily since relative tractor connections are compatible with tensorial constructions.
@@
So it remains to verify the claim about the representations corresponding to our relative tractor bundles, i.e. those obtained by tensorial constructions from the standard representation. Putting $\mathfrak{s}:=\mathfrak{sl}(n+1,\mathbb{R})\subset\mathfrak{p}/\mathfrak{p}_{+}$ semi-simplicity of $\mathfrak{s}$ implies that it acts trivially on any one-dimensional representation of $P/P_{+}$. Hence also the subgroup $SL(n+1,\mathbb{R})\subset P/P_{+}$ acts trivially so any such representation factorizes through the determinant representation $\det:GL(n+1,\mathbb{R})\to\mathbb{R}\setminus\{0\}$. The determinant representation is realized via the top exterior power of the standard representation. The square of this line bundle corresponds to a representation with values in $\mathbb{R}^{+}$, so we can form powers of that line bundle with arbitrary non-zero real exponents. These together with their tensor products with $\Lambda^{n}TM$ exhaust all possible on-dimensional representations of $P/P_{+}$.
@@
On the other hand, it is well known that any irreducible representation of $\mathfrak{s}$ can be realized in an iterated tensor product of copies of the standard representation and its dual via Young symmetrizers. By construction, any Young symmetrizer is $P/P_{+}$-equivariant, and hence each of these irreducible components is $P/P_{+}$-invariant. Thus we see that any irreducible representation of $\mathfrak{s}$ is realized by a tensorial construction and since forming direct sums is no problem, we can arrive at an isomorphic copy of any finite dimensional representation of $\mathfrak{s}$ from a tensorial construction. ∎
Remark 11.
@@
(1) It is clear that there are numerous example of Legendrean contact structures for which $F$ is involutive but the harmonic curvature component that is a section of the bundle $E^{*}\otimes F^{*}\otimes L(E,E)$ is non-zero. This follows for example from the relation of Legendrean contact structures to differential equations as used in [MFMZ]. In such cases, Theorem 10 applies to give a precise description of the sheaves resolved by relative BGG sequences. However, these geometries do not fall in the class of correspondence spaces as treated in part (2) of Theorem 4.11 of [Rel-BGG2]. Hence in these cases no explicit description of the resolved sheaves was known before.
@@
(2) The idea to use quotients of $TM$ as basic relative tractor bundles originally arose in the context of Legendrean contact structures during the work on the thesis [Michal-thesis] of the third author under the direction of the first author. Legendrean contact structures play only a minor role as an example in the thesis, however. The thesis mainly explores the fact that the machinery of relative BGG sequences extends beyond the case of Legendrean contact structures, namely to contact manifolds that are endowed with either a single involutive Legendrean distribution or with a Legendrean distribution with an arbitrary chosen complement in the contact distribution. In any case, the quotient of $TM$ by the distinguished Legendrean distribution plays a pivotal role in these developments, since this can be shown to carry a canonical partial linear connection obtained from the Bott connection as in §3.2 or from a family of distinguished partial linear connections determined by a choice of contact form, compare to Section 5.2.14 of [book] and to [Michal-article]. The latter are restrictions of Weyl connections, so one obtains descriptions of relative tractor bundles and relative tractor connections in terms of Weyl structures as discussed in Remark 6.
3.5. Example: Generalized path geometries
@@
These geometries represent a kind of a dual situation to the case of Legendrean contact structures discussed in §3.4 above, in the sense that locally they have a projectivized tangent bundle as an underlying structure. They can be defined directly by a configuration of distributions as follows. On a smooth manifold $M$ of dimension $12n+$ 1, one requires a distribution $H\subset TM$ of rank $n+1$ together with a decomposition $H=E\oplus V$ where $E$ and $V$ have rank $1$ and $n$, respectively, such that the bundle map $H\otimes H\to TM/H$ induced by the Lie bracket of vector fields vanishes on $V\otimes V$ and restricts to an isomorphism on $E\otimes V$.
@@
It then turns out that for $n\neq 2$, the distribution $V$ is involutive and locally $M$ is diffeomorphic to an open subset in the projectivized tangent bundle of a local leaf space for $V$. For $n=2$, we will assume that this is the case, since this is necessary in order to obtain relative BGG resolutions. Assuming that one deals with a global projectivized tangent bundle, $M=\mathcal{P}TN$, the subbundle $E\subset TM$ defines a $1$-dimensional foliation on $M$ and projecting the leaves to $N$, one obtains a family of paths (unparametrized curves) on $N$ with exactly one path through each point in each direction. This is a classical path geometry and generalized path geometries can be considered as a local analog of that (where, for example, paths can be defined for an open set of directions only). The importance of these structures comes from the fact that they provide a way to geometrize systems of second order ODE, which is the starting point for their study in [Fels].
@@
Similarly as in §3.4, the relation to parabolic geometries comes from the homogeneous model, which this time is the flag manifold $F_{1,2}(\mathbb{R}^{n+2})$ of lines in planes in $\mathbb{R}^{n+2}$. So we again put $G:=PGL(n+2,\mathbb{R})$ and the stabilizer $Q$ of $\mathbb{R}\subset\mathbb{R}^{2}\subset\mathbb{R}^{n+2}$ has the form $P\cap\tilde{P}$, where $G/P\cong\mathbb{R}P^{n+1}$ and $G/\tilde{P}$ is the Grassmannian $Gr(2,\mathbb{R}^{n+2})$. The distributions $V$ and $E$ are the vertical subbundles of the projections to $G/P$ and $G/\tilde{P}$, respectively, and the integral submanifolds of $E$ project exactly to the projective lines in $G/P$.
@@
On the level of Lie algebras, the situation is closely parallel to §3.4, in particular the bigrading induced by $\mathfrak{q}\subset\mathfrak{p}\subset\mathfrak{g}$ has the same form as in (3.2), but now with blocks of size $1$, $1$, and $n$, so now $\mathfrak{g}_{(-1,0)}$ is one-dimensional, while $\mathfrak{g}_{(0,-1)}$ and $\mathfrak{g}_{(-1,-1)}$ both have dimension $n$. In this case, the situation is less symmetric and we’ll make some remarks on the “other direction” $\mathfrak{q}\subset\tilde{\mathfrak{p}}\subset\mathfrak{g}$ below. We’ll also restrict to the case $n>1$ here, for $n=1$ there is no difference between generalized path geometries and Legendrean contact structures.
@@
The canonical Cartan geometry $(\mathcal{G}\to M,\omega)$ associated to $(M,H=E\oplus V)$ here has the property that the $Q$-invariant subspaces $(\mathfrak{g}_{(-1,0)}\oplus\mathfrak{q})/\mathfrak{q}$ and $(\mathfrak{g}_{(0,-1)}\oplus\mathfrak{q})/\mathfrak{q}$ of $\mathfrak{g}/\mathfrak{q}$ induce the subbundles $E$ and $V$ of $TM$, respectively. As in §3.4 one concludes that $T_{\rho}M=V$ and that the basic relative tractor bundle in this case is $\mathcal{V}M=TM/V$ with the natural filtration given by $H/V\subset TM/V$. Collecting our results in this case, we obtain:
Theorem 12.
@@
Let $(M,H=E\oplus V)$ be a generalized path geometry of dimension $2n+1\geq 5$ where $V$ is assumed to be involutive if $n=2$ and consider the relative tractor bundle $\mathcal{V}M:=TM/V$. Applying tensorial constructions to $\mathcal{V}M$ as in §2.3 we arrive at a class of relative tractor bundles that correspond to representations of $P/P_{+}\cong GL(n+1,\mathbb{R})$ which contain all one-dimensional representations and whose restrictions to $\mathfrak{sl}(n+1,\mathbb{R})\subset\mathfrak{p}/\mathfrak{p}_{+}$ exhaust all finite dimensional representations of this Lie algebra.
@@
For any of these relative tractor bundles and any local leaf space $\psi:U\to N$ for the distribution $F$ with connected fibers, the induced relative BGG sequence defines a fine resolution of the sheaf of pullbacks of sections of a tensor bundle on $N$. This tensor bundle is obtained by applying the tensorial constructions from above to the tangent bundle $TN$.
@@
According to Section 4.4.3 of [book], there are only two harmonic curvature components for generalized path geometries with involutive $V$. One of these is a torsion of homogeneity $2$ that is a section of $E^{*}\otimes(TM/H)^{*}\otimes V$. The other one is a curvature of homogeneity $3$ that is a section of $V^{*}\otimes(TM/H)^{*}\otimes L(V,V)$ (with additional symmetry properties). Using this information, the proof is completed exactly as for Theorem 10. ∎
Remark 13.

### v2

@@ 3.4 Example: Legendrean contact structures
Legendrean contact structures (also called Lagrangean contact structures following the article [20] in which they were first studied) are an example of parabolic contact structures, so they can be viewed as a refinement of contact structures. Explicitly, a Legendrean contact structure on a smooth manifold of odd dimension $12n+$ 1 is given by a contact structure $H\subset TM$ together with a decomposition $H=E\oplus F$ of $H$ as a direct sum of two subbundles of rank $n$, which are Legendrean in the sense that the Lie bracket of two sections of one of the subbundles always is a section of $H$. Otherwise put, for any $x\in M$ the fibers $E_{x},F_{x}\subset H_{x}\subset T_{x}M$ are maximal isotropic subspaces with respect to the non-degenerate bilinear form $\mathcal{L}_{x}\colon H_{x}\times H_{x}\to T_{x}M/H_{x}$ induced by the Lie bracket of vector fields. The interest in these structures mainly comes from their relation to the geometric theory of differential equations, see, e.g., [18] and to CR geometry, see, e.g., [12]. Indeed, Legendrean contact structures and partially integrable almost CR structures are two real forms of the same complex geometric structure.
@@
The relation to parabolic geometries comes from the homogeneous model of such geometries, which is the partial flag manifold $F_{1,n+1}\bigl{(}\mathbb{R}^{n+2}\bigr{)}$ of lines contained in hyperplanes in $\mathbb{R}^{n+2}$. This is a homogeneous space of the group $G:={\rm PGL}(n+2,\mathbb{R})$ and the stabilizer of the standard flag $\mathbb{R}\subset\mathbb{R}^{n+1}\subset\mathbb{R}^{n+2}$ descends to a parabolic subgroup $Q\subset G$. Indeed, $Q=P\cap\tilde{P}$ for two maximal parabolic subgroups of $G$, the stabilizers of $\mathbb{R}\subset\mathbb{R}^{n+2}$ and $\mathbb{R}^{n+1}\subset\mathbb{R}^{n+2}$. Correspondingly, there is a canonical homogeneous projection $G/Q\to G/P=\mathbb{R}P^{n+1}$ and it is easy to see that this identifies $G/Q$ with the projectivized cotangent bundle $\mathcal{P}\bigl{(}T^{*}\mathbb{R}P^{n+1}\bigr{)}$. This gives rise to a canonical contact structure on $G/Q$ for which the vertical subbundle is well known to be Legendrean. The second Legendrean subbundle making $G/Q$ into a Legendrean contact manifold is the vertical bundle of the canonical projection $G/Q\to G/\tilde{P}\cong\mathbb{R}P^{(n+1)*}$.
@@
$\begin{pmatrix}\mathfrak{g}_{(0,0)}&\mathfrak{g}_{(1,0)}&\mathfrak{g}_{(1,1)}\\ \mathfrak{g}_{(-1,0)}&\mathfrak{g}_{(0,0)}&\mathfrak{g}_{(0,1)}\\ \mathfrak{g}_{(-1,-1)}&\mathfrak{g}_{(0,-1)}&\mathfrak{g}_{(0,0)}\end{pmatrix}.$ (3.2)
@@
In particular, we just get one representation $\mathbb{V}_{i^{\prime}}$ in this case, namely $\mathbb{V}_{-1}=\mathfrak{g}/\mathfrak{p}$ and the $Q$-invariant filtration of this representation consists of the single invariant subspace $(\mathfrak{g}_{(-1,0)}\oplus\mathfrak{p})/\mathfrak{p}$. The corresponding bundles are easy to interpret, too.
@@
General results on parabolic geometries imply that for any Legendrean contact structure $(M,H=E\oplus F)$ there is a normal Cartan geometry $(p\colon\mathcal{G}\to M,\omega)$ of type $(G,Q)$ which is unique up to isomorphism, see [7, Section 4.2.3]. It is a general fact about Cartan geometries that $\mathcal{G}\times_{Q}(\mathfrak{g}/\mathfrak{q})\cong TM$ and the characterizing property of the geometry (apart from normality) is that the $Q$-invariant subspaces $\bigl{(}\mathfrak{g}_{(-1,0)}\oplus\mathfrak{q}\bigr{)}/\mathfrak{q}\subset\mathfrak{g}/\mathfrak{q}$ and $\bigl{(}\mathfrak{g}_{(0,-1)}\oplus\mathfrak{q}\bigr{)}/\mathfrak{q}\subset\mathfrak{g}/\mathfrak{q}$ correspond to the subbundles $E$ and $F$, respectively. Since $\mathfrak{g}_{(0,-1)}\oplus\mathfrak{q}=\mathfrak{p}$ we conclude that $T_{\rho}M=F$ in this case, and that our basic tractor bundle $\mathcal{V}_{-1}M$ is simply $TM/F$. The natural filtration of this bundle is given by the single invariant subspace $H/F\subset TM/F$.
@@
Now we can collect our results in the case of Legendrean contact structures. We restrict to the case $n>1$, i.e., $\dim(M)\geq 5$. The three-dimensional case can be easily dealt with directly. It is not very interesting from the point of view of relative BGG theory, however, since relative BGG sequences consist only of a single operator in this dimension.
@@
In the current situation, relative tractor bundles correspond to representations of the general linear group ${\rm GL}(n+1,\mathbb{R})$, so they correspond to natural vector bundles on smooth manifolds of dimension $n+1$. The relative tractor bundle $\mathcal{V}M=TM/F$ corresponds to the standard representation of ${\rm GL}(n+1,\mathbb{R})$ on $\mathbb{R}^{n+1}$, so via tensorial constructions as in Section 3.3 we obtain all analogs of tensor bundles. It is clear that the latter are by far the most important natural vector bundles (in fact, we are not aware of interesting applications of other types of natural bundles). Thus, it should be clear that we obtain (up to isomorphism) all the important relative tractor bundles on Legendrean contact structures via the constructions of Section 3.3, but it is difficult to make a precise statement in this direction. Instead of trying to do this, we just prove a result on the resulting representations in the following theorem, which shows the richness of the class.
Theorem 3.6.
@@
Let $(M,H=E\oplus F)$ be a Legendrean contact structure of dimension ${2n+1\geq 5}$ such that the distribution $F\subset TM$ is involutive and consider the relative tractor bundle $\mathcal{V}M:=TM/F$. Applying tensorial constructions to $\mathcal{V}M$ as in Section 3.3, we arrive at a class of relative tractor bundles that correspond to representations of $P/P_{+}\cong{\rm GL}(n+1,\mathbb{R})$ which contains all one-dimensional representations and whose restrictions to $\mathfrak{sl}(n+1,\mathbb{R})\subset\mathfrak{p}/\mathfrak{p}_{+}$ exhaust all finite-dimensional representations of this Lie algebra.
@@
For any of these relative tractor bundles and any local leaf space $\psi\colon U\to N$ for the distribution $F$ with connected fibers, the induced relative BGG sequence defines a fine resolution of the sheaf of pullbacks of sections of a tensor bundle on $N$. This tensor bundle is obtained by applying the tensorial constructions from above to the tangent bundle $TN$.
@@
We first need some information on the curvature of Legendrean contact structures that is available in the literature. For $n>1$, [7, Proposition 4.3.2] completely describes the harmonic curvature of Legendrean contact structures: There are two harmonic curvature quantities of homogeneity one, which are sections of the bundles $\Lambda^{2}E^{*}\otimes F$ and $\Lambda^{2}F^{*}\otimes E$, respectively. These turn out to be induced by the Lie bracket of vector fields and they are exactly the obstructions to involutivity of the distributions $E$ and $F$. So in particular, one of these components is exactly the obstruction to involutivity of $T_{\rho}M=F$. The third and last harmonic curvature quantity has homogeneity two and can be interpreted as a section of $E^{*}\otimes F^{*}\otimes L(E,E)$ (with some symmetries).
@@
Hence, involutivity of $F$ implies that the harmonic curvature vanishes upon insertion of two elements of $F$ while the components of torsion type even vanish upon insertion of one element of $F$. These are exactly the assumptions needed to apply part (1) of [10, Proposition 4.18]. The proof of this proposition then shows that the full Cartan curvature has the same properties, which in our language means that $\tau\bigl{(}T_{\rho}M,TM\bigr{)}=0$, so the assumptions of Corollary 3.5 are satisfied. On the other hand, [10, Proposition 4.18] says that the assumptions of part (1) of [10, Theorem 4.11] are satisfied. This shows that any relative BGG sequence is a fine resolution of the sheaf of local parallel sections for the relative tractor connection. Together with Corollary 3.5, this implies all claimed properties of the relative BGG sequence for $\mathcal{V}M=TM/F$. For bundles obtained via tensorial constructions, they then follow readily since relative tractor connections are compatible with tensorial constructions.
@@
So it remains to verify the claim about the representations corresponding to our relative tractor bundles, i.e., those obtained by tensorial constructions from the standard representation. Putting $\mathfrak{s}:=\mathfrak{sl}(n+1,\mathbb{R})\subset\mathfrak{p}/\mathfrak{p}_{+}$ semi-simplicity of $\mathfrak{s}$ implies that it acts trivially on any one-dimensional representation of $P/P_{+}$. Hence, also the subgroup ${\rm SL}(n+1,\mathbb{R})\subset P/P_{+}$ acts trivially so any such representation factorizes through the determinant representation $\det\colon{\rm GL}(n+1,\mathbb{R})\to\mathbb{R}\setminus\{0\}$. The determinant representation is realized via the top exterior power of the standard representation. The square of this line bundle corresponds to a representation with values in $\mathbb{R}^{+}$, so we can form powers of that line bundle with arbitrary non-zero real exponents. These together with their tensor products with $\Lambda^{n}TM$ exhaust all possible one-dimensional representations of $P/P_{+}$.
@@
On the other hand, it is well known that any irreducible representation of $\mathfrak{s}$ can be realized in an iterated tensor product of copies of the standard representation and its dual via Young symmetrizers. By construction, any Young symmetrizer is $P/P_{+}$-equivariant, and hence each of these irreducible components is $P/P_{+}$-invariant. Thus, we see that any irreducible representation of $\mathfrak{s}$ is realized by a tensorial construction and since forming direct sums is no problem, we can arrive at an isomorphic copy of any finite-dimensional representation of $\mathfrak{s}$ from a tensorial construction. ∎
Remark 3.7.
  1. (1)
@@
    It is clear that there are numerous examples of Legendrean contact structures for which $F$ is involutive but the harmonic curvature component that is a section of the bundle $E^{*}\otimes F^{*}\otimes L(E,E)$ is non-zero. This follows, for example, from the relation of Legendrean contact structures to differential equations as used in [18]. In such cases, Theorem 3.6 applies to give a precise description of the sheaves resolved by relative BGG sequences. However, these geometries do not fall in the class of correspondence spaces as treated in part (2) of [10, Theorem 4.11]. Hence, in these cases no explicit description of the resolved sheaves was known before.
  2. (2)
@@
    The idea to use quotients of $TM$ as basic relative tractor bundles originally arose in the context of Legendrean contact structures during the work on the thesis [22] of the third author under the direction of the first author. Legendrean contact structures play only a minor role as an example in the thesis, however. The thesis mainly explores the fact that the machinery of relative BGG sequences extends beyond the case of Legendrean contact structures, namely to contact manifolds that are endowed with either a single involutive Legendrean distribution or with a Legendrean distribution with an arbitrary chosen complement in the contact distribution. In any case, the quotient of $TM$ by the distinguished Legendrean distribution plays a pivotal role in these developments, since this can be shown to carry a canonical partial linear connection obtained from the Bott connection as in Section 3.2 or from a family of distinguished partial linear connections determined by a choice of contact form, compare to [7, Section 5.2.14] and to [21]. The latter are restrictions of Weyl connections, so one obtains descriptions of relative tractor bundles and relative tractor connections in terms of Weyl structures as discussed in Remark 3.2.
3.5 Example: Generalized path geometries
@@
These geometries represent a kind of a dual situation to the case of Legendrean contact structures discussed in Section 3.4 above, in the sense that locally they have a projectivized tangent bundle as an underlying structure. They can be defined directly by a configuration of distributions as follows. On a smooth manifold $M$ of dimension $12n+$ 1, one requires a distribution $H\subset TM$ of rank $n+1$ together with a decomposition $H=E\oplus V$ where $E$ and $V$ have rank $1$ and $n$, respectively, such that the bundle map $H\otimes H\to TM/H$ induced by the Lie bracket of vector fields vanishes on $V\otimes V$ and restricts to an isomorphism on $E\otimes V$.
@@
It then turns out that for $n\neq 2$, the distribution $V$ is involutive and locally $M$ is diffeomorphic to an open subset in the projectivized tangent bundle of a local leaf space for $V$. For $n=2$, we will assume that this is the case, since this is necessary in order to obtain relative BGG resolutions. Assuming that one deals with a global projectivized tangent bundle, $M=\mathcal{P}TN$, the subbundle $E\subset TM$ defines a one-dimensional foliation on $M$ and projecting the leaves to $N$, one obtains a family of paths (unparametrized curves) on $N$ with exactly one path through each point in each direction. This is a classical path geometry and generalized path geometries can be considered as a local analog of that (where, for example, paths can be defined for an open set of directions only). The importance of these structures comes from the fact that they provide a way to geometrize systems of second order ODE, which is the starting point for their study in [14].
@@
Similarly as in Section 3.4, the relation to parabolic geometries comes from the homogeneous model, which this time is the flag manifold $F_{1,2}\bigl{(}\mathbb{R}^{n+2}\bigr{)}$ of lines in planes in $\mathbb{R}^{n+2}$. So we again put $G:={\rm PGL}\bigl{(}n+2,\mathbb{R}\bigr{)}$ and the stabilizer $Q$ of $\mathbb{R}\subset\mathbb{R}^{2}\subset\mathbb{R}^{n+2}$ has the form $P\cap\tilde{P}$, where $G/P\cong\mathbb{R}P^{n+1}$ and $G/\tilde{P}$ is the Grassmannian ${\rm Gr}\bigl{(}2,\mathbb{R}^{n+2}\bigr{)}$. The distributions $V$ and $E$ are the vertical subbundles of the projections to $G/P$ and $G/\tilde{P}$, respectively, and the integral submanifolds of $E$ project exactly to the projective lines in $G/P$.
@@
On the level of Lie algebras, the situation is closely parallel to Section 3.4, in particular the bigrading induced by $\mathfrak{q}\subset\mathfrak{p}\subset\mathfrak{g}$ has the same form as in (3.2), but now with blocks of size $11$, and $n$, so now $\mathfrak{g}_{(-1,0)}$ is one-dimensional, while $\mathfrak{g}_{(0,-1)}$ and $\mathfrak{g}_{(-1,-1)}$ both have dimension $n$. In this case, the situation is less symmetric and we will make some remarks on the “other direction” $\mathfrak{q}\subset\tilde{\mathfrak{p}}\subset\mathfrak{g}$ below. We will also restrict to the case $n>1$ here, for $n=1$ there is no difference between generalized path geometries and Legendrean contact structures.
@@
The canonical Cartan geometry $(\mathcal{G}\to M,\omega)$ associated to $(M,H=E\oplus V)$ here has the property that the $Q$-invariant subspaces $\bigl{(}\mathfrak{g}_{(-1,0)}\oplus\mathfrak{q}\bigr{)}/\mathfrak{q}$ and $\bigl{(}\mathfrak{g}_{(0,-1)}\oplus\mathfrak{q}\bigr{)}/\mathfrak{q}$ of $\mathfrak{g}/\mathfrak{q}$ induce the subbundles $E$ and $V$ of $TM$, respectively. As in Section 3.4, one concludes that $T_{\rho}M=V$ and that the basic relative tractor bundle in this case is $\mathcal{V}M=TM/V$ with the natural filtration given by $H/V\subset TM/V$. Collecting our results in this case, we obtain.
Theorem 3.8.
@@
Let $(M,H=E\oplus V)$ be a generalized path geometry of dimension $2n+1\geq 5$, where $V$ is assumed to be involutive if $n=2$ and consider the relative tractor bundle $\mathcal{V}M:=TM/V$. Applying tensorial constructions to $\mathcal{V}M$ as in Section 3.3, we arrive at a class of relative tractor bundles that correspond to representations of $P/P_{+}\cong{\rm GL}(n+1,\mathbb{R})$ which contain all one-dimensional representations and whose restrictions to $\mathfrak{sl}(n+1,\mathbb{R})\subset\mathfrak{p}/\mathfrak{p}_{+}$ exhaust all finite-dimensional representations of this Lie algebra.
@@
For any of these relative tractor bundles and any local leaf space $\psi\colon U\to N$ for the distribution $F$ with connected fibers, the induced relative BGG sequence defines a fine resolution of the sheaf of pullbacks of sections of a tensor bundle on $N$. This tensor bundle is obtained by applying the tensorial constructions from above to the tangent bundle $TN$.
@@
According to [7, Section 4.4.3], there are only two harmonic curvature components for generalized path geometries with involutive $V$. One of these is a torsion of homogeneity $2$ that is a section of $E^{*}\otimes(TM/H)^{*}\otimes V$. The other one is a curvature of homogeneity $3$ that is a section of $V^{*}\otimes(TM/H)^{*}\otimes L(V,V)$ (with additional symmetry properties). Using this information, the proof is completed exactly as for Theorem 3.6. ∎
Remark 3.9.
  1. (1)